Biggers Inc., an all-equity firm is considering a change to its capital structure. The firm today is worth $750M. The firm’s management is considering borrowing $200M of debt and repurchasing their own stock with the money. The company historically has paid taxes at a rate of 30%; taxes are expected to remain at this level for the foreseeable future. a) IF the company expects to keep debt constant at $200M, what is the market cap, value of debt, and total levered value of the firm after the recapitalization? 1 b) IF the company intends to grow its debt by 3% per year AND the cost of the debt is 8%, what is the market cap, value of debt, and total levered value of the firm after the recapitalization? please show the detail step thx!
Answer #1

Value of all equity firm = $750m

Value of debt= 200(1-30%)= 140

Value of equity= 750-200= 550

Value of firm(leveraged)= 550+140=690